OpenStudy (anonymous):

slope 2, y-intercept -4 write this in an equation in either slope-intercept, point-intercept, or standard form.

OpenStudy (anonymous):

slope is "m", y-intercept is "b". slope-intercept form is y = mx + b. All you need to do now is make the substitutions.

OpenStudy (anonymous):

All the terms I used are in your question, so I'm confused about your question.

OpenStudy (anonymous):

Start with the general form of the equation, y = mx + b. That is slope-intercept. Just make the substitutions I identified for you.

OpenStudy (anonymous):

Are you still stuck?

OpenStudy (anonymous):

From your problem statement, slope is 2. I said slope is "m". So, "m" is 2. Also, from your problem statement, the y-intercept is -4. I said the y-intercept is "b". So, "b" is -4. Summarizing, "m" is 2 and "b" is -4. Equation is y = mx + b. So now, it should be easy to substitute into the equation now that you know m and b.